Find three consecutive whole numbers such that twice the sum of the two smallest numbers is 10 more than three times the largest number.
Answer by stanbon(75887)

Find three consecutive whole numbers such that twice the sum of the two smallest numbers is 10 more than three times the largest number.
----
1st: x-1
2nd: x
3rd: x+1
-----
Equation:
2(x-1+x)= 3(x+1)+10
4x-2 = 3x+13
x = 15
----
1st: x-1 = 14
2nd: 15
3rd 16
=================
Cheers,
Stan H.
==================